(1+2+3+4+5+6+7+8+9+0+1+2+3+4+5+6+7+8+9)-(1+2+3+4+5+6+7+8+9+0+1+2+3+4+5+6+7+8+9)

Grade:12th pass

1 Answers

YashRaj Soni
6 years ago
(47+0+47)-(47+47)=
94-94=0
-------------------------------------------------------------------------------------------------------------------------------------------------------------------============================-----------------------==========------------------------------==============-----------------------------------------------------------------------------------------------------------==============------------------------------------------======================------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------